OpenStudy (anonymous):

(2x)4 help ill give you a medal

OpenStudy (freckles):

(2x)4 or (2x)^4 ?

OpenStudy (freckles):

\[\text{ that is } \\ \text{ is it? } \\ (2x)4 \text{ or } (2x)^4 ?\]

OpenStudy (freckles):

\[(2x)4 \\ =(2 \cdot 4) x \]

OpenStudy (freckles):

\[(2x)^4 \\ =2^4 \cdot x^4\]